Define T : R3 → R2 by T(x,y,z) = (2x +4y +3z,6x)

Show that T is linear.
Answer #1

T: IR - IR by TA,7, 7) = (2x+4y 437, 6x) Now het d=(74,37971) ER, B =(13,95,22) tips. T()= T ( 64, 71, 71) = (224 +444+371